[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 14, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jul 14 15:21:00 UTC 2016


Here is today's summary of NHC computer operations:
--- Monica

----------

TAFB:
-Rainfall products haven't updated on website since Jul 7. TSB will  
investigate.
-The TAFB shift log scripts are intermittently reporting issues to the TSB  
spreadsheet. TSB is still investigating. Until that is fixed, TSB will have  
to start looking at the TAFB shift log responses spreadsheet since no email  
updates are being sent to TSB for TAFB technical issues.
HSU:
-Forecasters were having some ATCF GUI issues overnight. Rebooting the  
pac-atcf workstation seems to have corrected the problem.
-The adeck for Celia got corrupted overnight at sometime after 00Z leaving  
forecasters with no OFCLs for 00Z and 06Z cycles and a reduced amount of  
guidance for all of today. TSB is still troubleshooting to figure out the  
problem and working on re-creating a file that is as up to date as possible.
Other items:
-The networking problem within NWS yesterday (starting at 10:45 AM),  
prevented transmission of text products and loss of observations and SBN  
data was caused by a network engineer in Silver Spring forgetting to turn  
off a de-bugging process. Procedures are being put in place to prevent  
future occurrences. Fortunately this was right after the Celia and Darby  
products went out. Some TAFB products may have showed up as being late due  
to the outage. The connections were partially restored after about an hour,  
and fully restored after 4 hours.
-The Storm Surge AWIPS Sprint is continuing today
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160714/089718d5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list